A lab orders a shipment of 100 rats a week for experiments that the lab conducts. Prices for each weekly shipment of rats follow the distribution below. Suppose the mean cost of the rats turned out to be$ 13.00 per week. Interpret this value.

Price ​ $10.00 ​$12.50 ​$15.00

Probability. 0.25. 0.3. 0.45
Answer #1

Answer:

Given,

Mean = E(X) = x*P(x)

= 10*0.25 + 12.50*0.3 + 15*0.45

= 2.5 + 3.75 + 6.75

= 13

So the average cost for all weekly shipment of rats purchase is $ 13.